John was interviewed at 4 schools, the interview to acceptance percentage at the four schools were 45%, 35%, 70%, and 56%, what is johns chance of getting at least one acceptance to a school?
-
CycleTrack and SDN are teaming up to make medical school admissions more transparent and accessible! Read the announcement to learn how we’re supporting the future of CycleTrack’s free application-tracking tools .
You are using an out of date browser. It may not display this or other websites correctly.
You should upgrade or use an alternative browser.
You should upgrade or use an alternative browser.
Stats Question, need HELP please!!!
Started by iJackTeeth
Get help with your application
Use all the free resources available to you from SDN: articles, guides, expert advising, forums discussions, and school research.
Say the chance was 80% (it probably isn't). That means the chance he was accepted to NONE of the schools is 20%. So we want to solve for the chances he wasn't accepted anywhere and then figure out the chances he got at least 1 acceptance.John was interviewed at 4 schools, the interview to acceptance percentage at the four schools were 45%, 35%, 70%, and 56%, what is johns chance of getting at least one acceptance to a school?
P(no acceptance) = 55% * 65% * 30% * 44% (these are the percentages obtained when you do (100 - x) to all four percentages above - they represent the percentage of not getting accepted after an interview)
P(no acceptance) = (0.55)(0.65)(0.30)(0.44) = 0.04719
P(no acceptance) = 4.719%
P(acceptance somewhere) = 100 - 4.719 = 95.381%
Wats the answer? is it 51.5 percent ????
Say the chance was 80% (it probably isn't). That means the chance he was accepted to NONE of the schools is 20%. So we want to solve for the chances he wasn't accepted anywhere and then figure out the chances he got at least 1 acceptance.
P(no acceptance) = 55% * 65% * 30% * 44% (these are the percentages obtained when you do (100 - x) to all four percentages above - they represent the percentage of not getting accepted after an interview)
P(no acceptance) = (0.55)(0.65)(0.30)(0.44) = 0.04719
P(no acceptance) = 4.719%
P(acceptance somewhere) = 100 - 4.719 = 95.381%
oooohhhh,
now that makes more sense. thanks.Say the chance was 80% (it probably isn't). That means the chance he was accepted to NONE of the schools is 20%. So we want to solve for the chances he wasn't accepted anywhere and then figure out the chances he got at least 1 acceptance.
P(no acceptance) = 55% * 65% * 30% * 44% (these are the percentages obtained when you do (100 - x) to all four percentages above - they represent the percentage of not getting accepted after an interview)
P(no acceptance) = (0.55)(0.65)(0.30)(0.44) = 0.04719
P(no acceptance) = 4.719%
P(acceptance somewhere) = 100 - 4.719 = 95.381%
It makes so much sense when you explain in this. 🙄